OpenStudy (anonymous):

Mr. Garcia teaches French class. He has 14 students in his class, and the mean grade on his midterm was 72. After the midterm, one of his students, who had scored a 37 on the midterm, dropped out of the class. Mr. Garcia's recalculates the mean midterm grade, excluding the score from the student who dropped out. What is the recalculated mean midterm grade? Round your answer to the nearest tenth.

OpenStudy (anonymous):

mean of 14 students = 72 total of scores of 14 students = 14*72=1008 When one student goes number of students = 71 total of the scores of these 71 students = 1008-37 = 971 new midterm mean = 971/71 =13.676 = 13.7 rounded to the nearest tenth

OpenStudy (anonymous):

thank you thank you thank you

OpenStudy (anonymous):

welcome ☺